### v2

# Recent Advances in Online Hate Speech Moderation: Multimodality and the Role of Large Models

Abstract
In the evolving landscape of online communication, moderating hate speech (HS) presents an intricate challenge, compounded by the multimodal nature of digital content. This comprehensive survey delves into the recent strides in HS moderation, spotlighting the burgeoning role of large language models (LLMs) and large multimodal models (LMMs). Our exploration begins with a thorough analysis of current literature, revealing the nuanced interplay between textual, visual, and auditory elements in propagating HS. We uncover a notable trend towards integrating these modalities, primarily due to the complexity and subtlety with which HS is disseminated. A significant emphasis is placed on the advances facilitated by LLMs and LMMs, which have begun to redefine the boundaries of detection and moderation capabilities. We identify existing gaps in research, particularly in the context of underrepresented languages and cultures, and the need for solutions to handle low-resource settings. The survey concludes with a forward-looking perspective, outlining potential avenues for future research, including exploring novel AI methodologies, the ethical governance of AI in moderation, and developing more nuanced, context-aware systems. This comprehensive overview aims to catalyze further research and foster a collaborative effort towards more sophisticated, responsible, and human-centric approaches to HS moderation in the digital era.111 WARNING: This paper contains offensive examples.
**footnotetext: Both authors contributed equally to this work
Figure 1: Examples of an anti-migrant HS in different forms, encompassing text, image and/or audio modalities. The text-based, vision-language and video-based HS are taken from the Social Bias Inference Corpus (SBIC) dataset, the Facebook Hateful Memes (FHM) dataset and the Bitchute website, respectively.
1 Introduction
In the era of rapid information exchange and digital connectivity, the rise of hate speech (HS) presents a significant challenge with profound implications for global societies. Hate speech, which is any communication demeaning a person or group based on social or ethnic characteristics, undermines social harmony and individual safety, both online and offline Awan and Zempi (2016); Lupu et al. (2023); Wiedlitzka et al. (2023). The recent Israel-Hamas conflict has notably escalated both anti-Muslim and anti-Semitic sentiments worldwide, evidenced by the trending of hashtags such as #HitlerWasRight and #DeathToMuslim on social media platform X. [^2] https://www.nytimes.com/2023/11/15/technology/hate-speech-israel-gaza-internet.html Additionally, The Council on American-Islamic Relations reported receiving 774 help requests and bias reports from Muslims in the USA within a 16-day period. [^3] https://www.cair.com/press_releases/cair-reports-sharp-increase-in-complaints-reported-bias-incidents-since-107/ While digital interconnectivity facilitates swift information sharing, it simultaneously amplifies the spread and impact of HS, transcending geographical boundaries.
Technological advancements have transformed the expression of HS, leading to its manifestation in various novel forms. Traditionally, HS was predominantly text-based, found in written materials Ana (1999), or verbalized in posts, broadcasts, and public speeches Nielsen (2002). However, the digital era has ushered in more complex and subtle variants of HS, engaging multiple sensory modalities. A notable instance is vision-language HS, which fuses visual elements with textual content, commonly disseminated through captioned images and memes Kiela et al. (2020); Fersini et al. (2022); Bhandari et al. (2023). Video-based HS, another emerging form, amalgamates text, visuals, and audio, creating a multi-faceted and potentially more influential mode of communication Das et al. (2023). Figure 1 exemplifies various HS forms targeting immigrants, underscoring animosity towards individuals of diverse nationalities. The text-based approach overtly projects hostile attitudes towards immigrants in the host country. In vision-language HS, visual (e.g., a person preparing to shoot) and textual elements (e.g., sighting an illegal immigrant mowing the lawn) jointly convey antagonism. The figure also includes a music parody, integrating derogatory visuals with discriminatory audio lyrics, to showcase contempt for immigrants.
While the digital landscape has seen increasing diversity in HS forms, existing research surveys Rini et al. (2020); Chhabra and Vishwakarma (2023); Subramanian et al. (2023) predominantly concentrate on textual HS, often overlooking the complexity introduced by multimodal content. This survey aims to address this critical gap. Unlike prior surveys, we provide a comprehensive analysis of HS across various digital platforms, encompassing not only text but also visual, auditory, and combined multimodal expressions. We delve into the nuanced ways in which HS manifests in these different formats, offering insights into their unique characteristics and the challenges they pose for moderation.
Furthermore, our survey underscores the pivotal role of large language models (LLMs) and large multimodal models (LMMs) in moderating HS. These advanced models, which can process and interpret multiple types of data simultaneously, are crucial in understanding and countering the multifaceted nature of modern HS. We critically assess the existing solutions, evaluating their effectiveness and highlighting areas for improvement. Our survey presents a novel perspective on HS moderation, shifting the focus from traditional text-based analyses to multimodal approaches.
For this survey, we methodically examined research on moderating various types of hate speech, including text, images, videos, and audio. We searched with keywords like ‘hate speech’, ‘multimodal hate speech’, ‘harmful memes’, etc., on various scholarly platforms such as Google Scholar, DBLP, IEEE Xplore, and ACM Digital Library.
In summary, our paper not only bridges the gap in the existing literature by providing a detailed exploration of multimodal HS but also paves the way for future research in this area. We aim to inspire advancements in HS moderation technology, particularly in the development and refinement of large models, which are imperative for tackling the complex and ever-changing nature of online HS.
2 Hate Speech
Figure 2: Typology of HS based on modalities and tasks. The dark blue boxes are mature research areas with multiple studies; light grey boxes are ongoing research areas, while the hat can be found, and hatched boxes are unexplored topics.
HS takes various forms — written text, images, spoken words, and multimedia content — each posing risks of violence, animosity, or prejudice against specific groups. This section reviews existing literature on HS, categorizing it into text-based, image-based, video-based, and audio-based types. Figure 2 illustrates the range of online HS forms. Additionally, Table 1 lists publicly accessible HS datasets for different modalities, providing researchers with essential resources for data collection and analysis, and supporting comprehensive and consistent research in HS detection and moderation.
2.1 Text-based Hate Speech
Text-based HS encompasses written or typed expressions manifesting across various digital formats, including social media posts Waseem and Hovy (2016); Founta et al. (2018). Researchers addressing online HS have compiled datasets to detect and analyze its complexity and multifaceted nature, as detailed in existing literature. These studies explore diverse aspects of hateful and derogatory language, focusing on themes such as implicit HS ElSherief et al. (2021), targeted groups Kennedy et al. (2018), and types of attacks ElSherief et al. (2021). As detection models evolve, it becomes imperative to understand and elucidate their decision-making processes, leading to the development of datasets for implicit statements and methodologies for bias analysis Sap et al. (2020); ElSherief et al. (2021). Additionally, some research shifted towards proactive strategies, including countering HS Masud et al. (2022).
The detection of text-based HS is fraught with significant challenges. A primary obstacle is contextual understanding; HS often subtly or covertly conveys its message through sarcasm, irony, or cultural references. Linguistic variations, including slang, dialects, and unconventional language forms, exacerbate this intricacy, posing difficulties for automated recognition systems. Additionally, the multilingual nature of online content necessitates language-specific tools or models for effective detection across diverse linguistic contexts. A further complication is the rapid evolution of online language, which can quickly render existing detection models outdated.
2.2 Image-based Hate Speech
Image-based HS utilizes visual elements, such as photographs, cartoons, and illustrations, to propagate hate or discrimination against specific groups. A common manifestation of this type is memes, which typically consist of images combined with short overlaid text. Although memes often serve humorous or satirical purposes, they are increasingly used to spread hateful content online Kiela et al. (2020). The viral nature of memes on digital platforms further amplifies their impact, allowing them to reach vast audiences rapidly. To address the proliferation of hateful memes, recent studies have focused on developing datasets for their detection Kiela et al. (2020); Gomez et al. (2020) and for identifying specific targets and types of attacks within these memes Fersini et al. (2022). Beyond detection, new approaches are being pursued to analyze and mitigate bias in image-based HS detection models Hee et al. (2023); Lin et al. (2023). Additionally, new methodologies are emerging to counteract HS transmitted through memes Van and Wu (2023).
Detecting image-based HS presents challenges due to the nuanced expression of offensive messages in visual content. Images, often embedding symbols, memes, or culturally specific visual cues, require deep cultural and contextual understanding for accurate interpretation. The combination of visual elements and text can subtly imply meanings not immediately evident Kiela et al. (2020). For example, Figure 1 depicts a man with a gun and text suggesting hostility towards immigrants. Differentiating humor from hate in memes is particularly challenging, influenced by varying cultural, societal, and personal perspectives Schmid (2023).
2.3 Video-based Hate Speech
Video-based HS presents a complex challenge, comprising a blend of visuals, audio tracks, and/or textual elements. This form of HS ranges from professionally produced propaganda to amateur videos on social media platforms like YouTube and TikTok. The engaging nature of video content and its easy dissemination across digital networks significantly heighten its potential for harm. Echoing the concerns of image-based HS, video-based HS also contributes to the normalization of hateful ideologies and can profoundly influence public opinion. Contemporary research in this domain primarily focuses on identifying video-based HS and categorizing its various subtypes Wu and Bhandary (2020). Recent studies have explored detecting hateful segments within videos alongside multi-dimensional analysis. Nonetheless, the body of research on video-based HS is less developed than text-based and image-based HS, particularly in areas such as analyzing and mitigating model bias, elucidating decision-making processes, and devising counterstrategies. These gaps, likely stemming from the rapid pace of technological advancements and evolving digital trends, underscore the need for further research to promote a more harmonious online environment.
Identifying HS within videos is a demanding and resource-intensive endeavor. This complexity stems from the amalgamation of various elements, such as text, images, and audio, within a single medium. Each component can independently harbor HS, thereby compounding the detection process. The duration of videos further exacerbates this challenge, as longer content necessitates more extensive review and analysis, with potential shifts in context over time. Moreover, subtle visual cues and sophisticated editing techniques can be employed to discreetly embed hate messages, making their detection by automated tools particularly challenging. Additionally, the analysis of video content requires considerable computational resources and time, posing a substantial challenge for organizations tasked with detecting and addressing HS in video formats.
2.4 Audio-based Hate Speech
Audio-based HS entails the analysis of sound waves to discern elements such as pitch, intonation, and the contextual meaning of spoken words. This form of HS can originate from a variety of audio channels, including real-time conversations, podcasts, broadcasts, and other forms of audio media. The methodologies for addressing audio-based HS are diverse, targeting different facets of the issue. For instance, Barakat et al. Barakat et al. (2012) employed a straightforward keyword-based approach to identify segments of HS, while Wazir et al. Wazir et al. (2020) engaged in a detailed classification of offensive categories in audio-based HS, showcasing a nuanced method of understanding and categorizing this form of HS. This research area is still in its developmental stages, partly due to the scarcity of dataset. Nonetheless, recognizing the variety and significance of the approaches and techniques employed in this field is imperative. This recognition not only sheds light on the current state of research but also illuminates potential avenues for future exploration.
Detecting HS in audio recordings presents unique challenges, primarily related to the transcription and interpretation of spoken words. The accuracy of speech recognition is crucial, especially when dealing with diverse accents, background noise, or poor audio quality. Additionally, the tone and intonation of spoken language play a significant role in conveying intent, which can substantially alter the meaning of words. This aspect poses a challenge for detection based solely on text transcripts, as subtle nuances in vocal expression may be lost during transcription. Moreover, non-verbal audio elements, such as sound cues or background noises, are pivotal in contextualizing speech. However, these elements are often difficult to interpret using automated methods.
3 Methodology
Mod. Dataset Task Labels Source # Records
Text WZ-LS Waseem and Hovy (2016) Det.
[M.C.] Sexism, Racism, Neither
Twitter 16,914
GHC Kennedy et al. (2018) Det.
[M.C.] VO, HD, CV [B] Implicit, Explicit [M.C.] Hate Targets
Forums 27,665
Stormfront de Gibert et al. (2018) Det.
[B] Hateful
StormFront 9,916
DT Davidson et al. (2017) Det.
[M.C.] Hateful, Offensive, Neither
Twitter 24,802
HatEval Basile et al. (2019) Det.
[B] Hateful [B] Target [B] Aggressive
Twitter 19,600
Founta Founta et al. (2018) Det.
[M.C.] Offensive, Abusive, Hateful Speech, Aggressive, Cyberbullying, Spam, Normal
Twitter 80,000
SBIC Sap et al. (2020) Det.
[B] Offensive [B] Intent [B] Lewd [B] Group [B] Hate Targets [B] In-Group
Mixed 44,671
Expl.
[F.T.] Implied Statement
IHC ElSherief et al. (2021) Det.
[M.C.] Implicit, Explicit, Non-Hate [M.C.] Grievance, Incitement, Inferiority, Irony, Stereotypical, Threatening, Others
Twitter 22,584
Expl.
[F.T.] Implied Statement
HateXplain Mathew et al. (2021) Det.
[M.C] Hate, Offensive, Normal [M.C.] Hate Targets
Mixed 20,148
Expl.
[M.L] Text Rationales/Snippets
DynaHate Vidgen et al. (2021) Det.
[B] Hateful, [M.C] Animosity, Derogation, Dehumanization, Threatening, Support, [M.C.] Hate Targets
H-M Adv 41,134
ToxiGen Mathew et al. (2021) Det.
[B] Toxic, Benign
GPT-3 274,186
NACL Masud et al. (2022) Det.
[M.C.] Hate Intensity [M.L.] Hate Spans
Mixed 4,423
Ctr.
[F.T.] Hate Speech Normalization
CONAN Chung et al. (2019) Det.
[M.C.] Hate Types [M.C.] Hate Sub-Topic
Synthetic 14,988
Ctr.
[F.T.] CN Generation
Multitarget CONAN Fanton et al. (2021) Det.
[M.C.] Hate Targets
GPT-2 5,003
Ctr.
[F.T.] CN Generation
Counter Narratives Das et al. (2023) Ctr.
[F.T.] CN Generation
YouTube 9,119
Img MMHS150K Gomez et al. (2020) Det.
[B] Hateful
Twitter 150,000
FHM Kiela et al. (2020) Det.
[B] Hateful
Synthetic 10,000
Finegrained FHM Mathias et al. (2021) Det.
[B] Hateful [M.L.M.C] Protected Category [M.L.M.C] Protected Attacks
Synthetic 10,000
Misogynous Meme Det.
[B] Misogynistic [B] Aggressive
Mixed 800
Gasparini et al. (2022)
[B] Ironic
MAMI Fersini et al. (2022) Det.
[B] Misogyny [M.L.M.C.] Misogynous, Shaming, Stereotype, Objectification, Violence
Mixed 10,000
UA-RU Conflict Thapa et al. (2022) Det.
[B] Hateful
Twitter 5,680
CrisisHateMM Bhandari et al. (2023) Det.
[B] Hateful [B] Directed [M.C.] Hate Targets
Mixed 4,723
HatReD Hee et al. (2023) Expl.
[F.T] Explanations
Synthetic 3,228
Video HateMM Das et al. (2023) Det.
[B] Hateful [M.C.] Hate Targets
Mixed 1,083
Bangla Hate Videos Junaid et al. (2021) Det.
[B] Hateful
YouTube 300
Audio DeToxy Ghosh et al. (2021) Det.
[B] Hateful / Non-hateful
Mixed 2M
MuTox Costa-jussà et al. (2024) Det.
[B] Hateful / Non-hateful
Mixed 116,000
Table 1: Publicly available datasets for HS detection (Det.), HS explanation (Expl.) and counter HS (Ctr.). Abbreviation: M.L.: multi-label, M.C.: multi-class, M.L.M.C.: multi-label multi-class, B: binary, F.T: free-text, H-M Adv: Human-Machine Adversarial. Note that multilingual HS is out of the scope for the current review.
In this section, we review state-of-the-art methodologies that have made significant contributions to key areas of HS research: HS detection, HS explanation, HS debiasing, and counter speech. We emphasise research studies incorporating LLMs or LMMs, aiming to garner comprehensive insights into emerging trends within this domain. This allows us to understand how these advanced models enhance our understanding and management of HS in various forms.
3.1 Large Models
The emergence of large foundation models, such as LLMs and LMMs, marks a significant milestone in artificial intelligence research, showcasing unprecedented capabilities in understanding and generating data across different formats Zhao et al. (2023). LLMs are designed to excel in language understanding and text generation Touvron et al. (2023). In contrast, LMMs are adept at processing and interpreting various data types, including visual, textual, and auditory inputs, enabling a broader spectrum of applications Yang et al. (2023b). These foundation models have opened new avenues for identifying and mitigating hateful content, which requires nuanced understanding of language and context.
In this survey, we regard both LLMs and LMMs as models with several billion parameters, aligning with the definition widely accepted and analyzed in numerous studies of large-scale models Luo et al. (2023).
3.2 Hate Speech Detection
The leading detection techniques for HS vary according to the modality of the content, encompassing approaches from transformer-based models to spectrogram-based classification models. For instance, AngryBERT Awal et al. (2021) employs a multi-task learning approach to fine-tune BERT specifically for binary text-based HS detection. PromptHate Cao et al. (2022) integrates in-context learning with a demonstration sampling strategy for mask prompt fine-tuning on RoBERTa, targeting hateful memes. In audio-based HS classification, studies Boishakhi et al. (2021); Ibañez et al. (2021) have leveraged ensemble techniques involving AdaBoost, Naive Bayes, and random forest. Additionally, CNNs have been utilized to convert audio into spectrograms Medina et al. (2022), with self-attentive CNNs Yousefi and Emmanouilidou (2021) employed to extract audio features. For video-based HS classification, a combination of BERT, ViT, and MFCC has been used for text, image, and audio modality analysis, respectively Das et al. (2023). Note that video-based and audio-based HS detection represent emerging fields, offering substantial scope for further development and innovation.
Transformer-based models have significantly advanced the detection of text-based and image-based HS; yet they encounter specific challenges. For text-based models, a major hurdle is generalizing to out-of-distribution datasets, often hindered by limited vocabulary and the rarity of implicit HS in many datasets Ocampo et al. (2023b). To overcome this, recent initiatives include adversarial HS generation and in-context learning with LLMs. Ocampo et al. Ocampo et al. (2023a) introduced a method using GPT-3 to generate implicit HS, aiming to both challenge and improve HS classifiers. Concurrently, Wang et al. Wang et al. (2023b) developed a technique for optimizing example selection for in-context learning in LLMs.
In image-based HS, the primary challenge lies in deciphering implicit hate messages within memes. This difficulty often stems from the loss of information during the extraction of text-based features from images, a common step in many methodologies Lee et al. (2021); Pramanick et al. (2021); Cao et al. (2022). Furthermore, the implicit HS in memes can be concealed by seemingly unrelated text and images, as illustrated in Figure 1. To address these challenges, recent strategies include employing LMMs with prompting techniques and/or knowledge distillation. ProCap Cao et al. (2023) addresses the issue of information loss in image-to-text conversion by prompting an LMM in a QA format, enhancing the generated caption’s quality and informativeness. To tackle the problem of disconnected text and images, MR.HARM Lin et al. (2023) utilizes an LMM to generate potential rationales. These rationales are subsequently employed to fine-tune supervised HS classification systems through knowledge distillation, improving the detection of hateful memes.
3.3 Hate Speech Explanation
A major challenge in contemporary HS detection methods is their lack of explainability in decision-making processes. Explainability is crucial for fostering user trust and facilitating systems that require human interaction Balkir et al. (2022). One proposed solution involves training supervised models that not only categorize HS but also provide rationales for these classifications Sap et al. (2020); ElSherief et al. (2021); Hee et al. (2023). Sap et al. Sap et al. (2020) and Elsherief et al. ElSherief et al. (2021) developed text-based HS datasets with human-annotated explanations, setting benchmarks for identifying underlying hate. Similarly, Hee et al. Hee et al. (2023) compiled a dataset for hateful memes, complete with human-annotated explanations and benchmarks. However, the process of gathering human-written explanations is not only time-consuming but also susceptible to individual biases. Moreover, it involves the risk of subjecting human annotators to prolonged exposure to HS, which can have adverse psychological effects.
Recent studies have delved into employing LLMs to generate plausible and meaningful explanations for HS. For instance, Wang et al. Wang et al. (2023a) demonstrated that GPT-3 can craft convincing and effective explanations for HS, a finding substantiated by extensive human evaluations. Additionally, HARE Yang et al. (2023a) introduces two innovative methods for generating rationales in HS detection, which have been shown to improve both the training process and the performance of detection models. This approach presents an alternative means of developing insightful explanations, while simultaneously mitigating the risks associated with prolonged human exposure to HS. Nevertheless, this area of research is still nascent, thus presenting numerous opportunities for further investigation and development.
3.4 Hate Speech Debiasing
Bias in HS detection models poses a significant risk to their effectiveness and fairness, leading to potential adverse impacts on individuals and society. Addressing this, numerous studies have focused on identifying and mitigating bias in these models. Sap et al. Sap et al. (2019) found that two widely-used corpora exhibit bias against African American English, which increases the likelihood of classifying tweets in this dialect as hateful. Hee et al. Hee et al. (2022) conducted a quantitative analysis of modality bias in hateful meme detection, observing that the image modality significantly influences model predictions. Their study also highlighted the tendency of these models to generate false positives when encountering specific group identifier terms.
Beyond merely identifying biases, various studies have introduced innovative methods to reduce these biases within models. Kennedy Kennedy et al. (2020) developed a regularization technique utilizing SOC post-hoc explanations to address group identifier bias. Similarly, Rizzi et al. Rizzi et al. (2023) observed that models exhibit biases towards terms linked with stereotypical notions about women, such as dishwasher and broom. To counteract this, the authors proposed a bias mitigation strategy using Bayesian Optimization, which effectively lessened the bias while preserving overall model performance.
These efforts underscore the critical importance of not only recognizing but actively mitigating bias in HS detection models. This is especially vital as large models increasingly dominate the landscape for generating explanations and enabling transfer learning. Addressing biases within these large models represents a significant opportunity to advance new methodologies and strategies for both the identification and reduction of bias.
3.5 Counter Speech
The approach to counter HS emphasizes creating non-aggressive reactions, aimed at curbing the spread of HS and potentially converting it into respectful and inoffensive language. The primary approach for counter HS involves producing counter-narratives (CN), which rely on reliable evidence, logical arguments, and different perspectives to challenge the hate directed at marginalized groups Chung et al. (2021). Multitarget CONAN Fanton et al. (2021) uses GPT-2 to generate potential CN candidates and then engages human evaluators to identify effective CNs, leading to a dataset of 4,07840784,0784 , 078 CNs. An alternative approach involves either diminishing (i.e., normalization) or eliminating (i.e., correction) the level of hate in HS. NACL Masud et al. (2022) employs neural networks to paraphrase sections of text containing hate, effectively lessening the intensity of hate in HS. Van and Wu Van and Wu (2023) utilized LMM to identify and correct HS, substituting the text in memes with positive and respectful language.
These studies underscore the critical role of generative models in annotating and developing counter speech strategies. This further signifies the future opportunities of LMMs and LMMs in enhancing approaches and techniques to combat hate speech effectively.
4 Challenges
In the dynamic realm of research, especially in areas related to user-generated content and online harmfulness, numerous challenges persist that shape the trajectory and emphasis of scholarly investigations. These challenges, ranging from technical to ethical, define the landscape in which research on HS moderation and detection operates.
Data Complexity, Quality, and Sourcing.
Addressing multimodal HS brings to the fore complex challenges concerning data complexity, quality, and sourcing. In text and image-based contexts, issues range from reliance on pre-trained models to the nuanced interpretation of visual cues Cao et al. (2020); Awal et al. (2021); Sridhar and Yang (2022). For audio and video content, challenges are compounded by factors like varied accents, background noise, and the ambiguous nature of toxic content, coupled with inconsistent audio quality Yousefi and Emmanouilidou (2021). Furthermore, sourcing data from diverse platforms such as Gab, YouTube, and 4chan introduces difficulties in standardization and interpretation Mariconti et al. (2019). Additionally, the uneven distribution of profanity across datasets poses significant obstacles for accurate model training. These challenges underscore the need for advanced methods capable of navigating the intricate and multifaceted nature of data across different modalities.
Model Performance and Generalizability.
Recent research highlights the importance of enhancing HS detection models for adaptability in various scenarios and contexts. An exemplary example is making HS detection generalizable and effective across different domains Awal et al. (2021), underscoring the need for models to be versatile and not overly reliant on specific content cues such as domain, region, demography, and more. The development of systems like VulnerCheck Mariconti et al. (2019) exemplifies the demand for models that perform well regardless of the context, emphasizing the importance of creating models that can adapt to the ever-evolving nature of online material. Such adaptability is crucial for identifying and managing new hateful content, especially those designed to bypass advanced AI technologies. The adoption of technologies, like Few-Shot Learner (FSL), for quick adaptation to this evolving landscape shows a promising direction [^4] https://ai.meta.com/blog/harmful-content-can-evolve-quickly-our-new-ai-system-adapts-to-tackle-it/. However, it is imperative that these technologies not only understand the content but also integrate critical aspects of cultural, behavioral, and conversational contexts into their analysis.
Modality and Expression Variabilities.
Research has highlighted the complexities involved in interpreting expressions of HS across individual and combination of various modalities, particularly in cyberbullying scenarios Boishakhi et al. (2021). The main challenges in this area include the limitations of pre-trained models in accurately capturing the subtleties of HS expressed via multiple modalities Cao and Lee (2020), the difficulty in balancing sensitivity and specificity in detection algorithms Cao et al. (2020), and the imperative for models to distinguish between explicit and implicit forms of HS Kim et al. (2022). Additionally, the potential for misinterpretation and the sensitivity of models to specific trigger language Sridhar and Yang (2022) pose significant challenges. Addressing these issues is crucial and requires concerted efforts in future research endeavors.
Contextualization.
Recent research underscores the complexities of identifying and countering online hate speech (HS), requiring an accurate and comprehensive understanding of HS, ranging from the targeted victims to its implicit meanings Sap et al. (2020); Masud et al. (2022). An effective strategy to counter hate speech depends on this precise interpretation, enabling the creation of context-specific responses. The main obstacles in accurately decoding hate speech involve recognizing the subtle nuances in language use, such as wordplays, dual meanings, or ambiguous expressions Kennedy et al. (2020). Additionally, the interpretation of hate speech often depends on its context, including cultural, societal, and situational elements Rizzi et al. (2023). These challenges highlight the need for sophisticated algorithms capable of interpreting language within its contextual usage, thereby enhancing the accuracy and effectiveness of hate speech detection and mitigation strategies.
Emerging Domains.
Exploring new and evolving fields, such as the metaverse, presents a distinct set of challenges Medina et al. (2022). The core of these challenges lies the need to adapt current hate speech (HS) detection methods to new contexts and to develop new strategies specifically designed for the unique characteristics of these platforms. The dynamic and immersive nature of these emerging environments necessitates a re-evaluation and potential re-engineering of current HS detection and mitigation strategies. Future research requires a deep understanding of both technological advancements and the social dynamics within these virtual spaces to ensure effectiveness in detecting and mitigating HS within these evolving digital landscapes.
Bias and Ethical Concerns:
Confronting bias and upholding ethical considerations in AI systems, particularly in the context of HS detection, represents a challenge, as detailed in works such as Sandulescu (2020). These issues are not just technical but also moral, encompassing the need to ensure that AI systems operate equitably and do not inadvertently perpetuate or amplify existing societal biases. The development and implementation of responsible AI systems, therefore, require a multidisciplinary approach that integrates technical proficiency with ethical and societal awareness. This necessity underlines the importance of continuous efforts to refine AI algorithms, ensuring they align with ethical standards and societal values.
In summary, the areas of HS detection and moderation are confronted with multifaceted challenges. These arise from the inherent complexities of data, technological limitations, modality variabilities, dataset biases, and the uncharted territories of emerging domains like the metaverse. To effectively navigate these obstacles, a concerted and multidisciplinary effort is essential. It calls for the development of methodologies that are not only sophisticated and robust but also highly adaptable. Such methodologies must be capable of contending with the dynamic and often unpredictable nature of user-generated content and online interactions. The future of this field hinges on our ability to continuously evolve and innovate, ensuring that our approaches remain relevant, effective, and ethically sound in an ever-changing digital landscape.
5 Future Directions: Role of Large Models
To conclude this survey, we explore prospective research opportunities in utilizing large-scale models for moderating hate content.
Cross-Modality Context Understanding.
As hate speech extends beyond mere text to encompass multiple forms of media (multimodality), it becomes crucial for models to possess a proficient understanding of context across these different modalities. Hence, it is imperative that models not only identify hateful content within text or images separately but also grasp how the combination of text and images can alter the overall message Kiela et al. (2020). For instance, an image that is benign on its own might become hateful when paired with specific text. Research could focus on developing models that more effectively understand context across different modalities.
Low-Resource Hate Speech Adaptation.
Domain adaptation between related tasks has gained significant attention, with transfer learning emerging as a prominent strategy for addressing this challenge. In the domain of hate speech, an exemplary application is the cross-lingual transfer learning for detecting hate speech across different languages. Winata et al. Winata et al. (2022) use few-shot in-context learning and fine-tuning techniques to adapt insights from languages with abundant resources to those with fewer resources. Given the widespread presence of hate speech and its relatively consistent definitions across different forms, there is potential to extend knowledge from text-based hate speech with abundant resources to other low-resource forms of hate speech. Future research should aim to develop models capable of pre-training on a broad spectrum of multimodal data, including text, images, audio, and videos, to enhance transfer learning capabilities.
Humour & Sarcasm Understanding.
Comprehending humor and sarcasm involves recognizing subtle linguistic signals and understanding the broader context, which includes cultural, social, and environmental factors. LLMs are adept at processing language but might not entirely capture these intricate details or fully understand the specific circumstances surrounding a statement. Additionally, humor and sarcasm often hinge on wordplay, double meanings, or ambiguous interpretations. Although LLMs can identify language patterns, they might struggle to differentiate between straightforward and figurative speech. Research efforts can focus on enhancing the capability of LLMs and LMMs to interpret sarcasm and humor, particularly dark humor, which conceals itself within the context of a sentence.
Multicultural Moderation.
A challenge in hate speech detection lies in the varying cultural and contextual cues across different countries and regions. These subtle cues often require a nuanced understanding of local languages, dialects, slang, and social norms. This complexity makes it difficult for automated systems to identify and differentiate hate speech from non-offensive content. Nguyen et al. Nguyen et al. (2023) demonstrated how providing cultural common-sense knowledge can alter GPT-3’s behaviour, leading it to produce more accurate and culturally sensitive questions. Similarly, future research could aim to curate HS dataset with regional culture information and build culturally-aware LLMs and LMMs by injecting and fine-tuning models with cultural knowledge.
Real-Time Monitoring.
The vocabulary of hate speech is constantly changing and evolving, particularly in online spaces. Although adapting to different domains can enhance a model’s capacity to apply its knowledge across a range of current datasets, the ongoing development of new slurs, coded terms, and symbolic expressions presents a considerable obstacle to the successful detection of hate speech. Research efforts can focus on continual learning methods that enable these models to be updated regularly while minimising the adjustments to their parameters.
Factual Grounding.
Although current methods in generating HS explanation using large-scale models (refer to Section 3.3) have shown promise, they still face significant challenges. These large models are prone to “hallucinations” producing responses that can be factually incorrect, illogical, or unrelated to the initial prompt Ji et al. (2023). Consequently, while these recent advancements are promising, the explanations generated by these models are susceptible to misinformation and require verification. Future research should aim to improve the accuracy and relevance of these explanations, which might involve anchoring the explanations in verifiable facts and developing techniques to identify and rectify any discrepancies.
6 Conclusion
This survey has highlighted significant advancements in HS moderation, underscoring the pivotal role of large language and multimodal models. Despite these strides, challenges remain, particularly in inclusivity and nuanced detection. Future research should focus on developing AI methodologies that are more context-aware and ethically governed. This endeavor is not only a technological challenge but also a moral imperative, necessitating interdisciplinary collaboration. As we advance, it is crucial to balance innovation with responsibility, striving for a digital landscape that is safer and more inclusive for all.

### v4

@@
 Abstract
-In the evolving landscape of online communication, moderating hate speech (HS) presents an intricate challenge, compounded by the multimodal nature of digital content.
-This comprehensive survey delves into the recent strides in HS moderation, spotlighting the burgeoning role of large language models (LLMs) and large multimodal models (LMMs).
-Our exploration begins with a thorough analysis of current literature, revealing the nuanced interplay between textual, visual, and auditory elements in propagating HS.
-We uncover a notable trend towards integrating these modalities, primarily due to the complexity and subtlety with which HS is disseminated.
-A significant emphasis is placed on the advances facilitated by LLMs and LMMs, which have begun to redefine the boundaries of detection and moderation capabilities.
-We identify existing gaps in research, particularly in the context of underrepresented languages and cultures, and the need for solutions to handle low-resource settings.
-The survey concludes with a forward-looking perspective, outlining potential avenues for future research, including exploring novel AI methodologies, the ethical governance of AI in moderation, and developing more nuanced, context-aware systems.
-This comprehensive overview aims to catalyze further research and foster a collaborative effort towards more sophisticated, responsible, and human-centric approaches to HS moderation in the digital era.111 WARNING: This paper contains offensive examples.
-**footnotetext: Both authors contributed equally to this work
+Moderating hate speech (HS) in the evolving online landscape is a complex challenge, compounded by the multimodal nature of digital content.
+This survey examines recent advancements in HS moderation, focusing on the burgeoning role of large language models (LLMs) and large multimodal models (LMMs) in detecting, explaining, debiasing, and countering HS.
+We begin with a comprehensive analysis of current literature, uncovering how text, images, and audio interact to spread HS.
+The combination of these modalities adds complexity and subtlety to HS dissemination.
+We also identified research gaps, particularly in underrepresented languages and cultures, and highlight the need for solutions in low-resource settings.
+The survey concludes with future research directions, including novel AI methodologies, ethical AI governance, and the development of context-aware systems.
+This overview aims to inspire further research and foster collaboration towards responsible and human-centric approaches to HS moderation in the digital age.111 WARNING: This paper contains offensive examples.
+Recent Advances in Online Hate Speech Moderation:
+Multimodality and the Role of Large Models
+Ming Shan Hee1thanks: These authors contributed equally to this work., Shivam Sharma2,*, Rui Cao3, Palash Nandi2, Preslav Nakov4, Tanmoy Chakraborty2, Roy Ka-Wei Lee1, 1SUTD, 2IIT Delhi, 3SMU, 4MBZUAI 1{mingshan_hee@mymail., roy_lee@}sutd.edu.sg 2{shivam.sharma, palash.nandi, tanchak}@ee.iitd.ac.in 3ruicao.2020@phdcs.smu.edu.sg 4preslav.nakov@mbzuai.ac.ae
 Figure 1: Examples of an anti-migrant HS in different forms, encompassing text, image and/or audio modalities.
@@
 In the era of rapid information exchange and digital connectivity, the rise of hate speech (HS) presents a significant challenge with profound implications for global societies.
-Hate speech, which is any communication demeaning a person or group based on social or ethnic characteristics, undermines social harmony and individual safety, both online and offline Awan and Zempi (2016); Lupu et al.
-(2023); Wiedlitzka et al.
+HS, which is any communication demeaning a person or a group based on social or ethnic characteristics, undermines social harmony and individual safety, both online and offline Lupu et al.
 (2023).
-The recent Israel-Hamas conflict has notably escalated both anti-Muslim and anti-Semitic sentiments worldwide, evidenced by the trending of hashtags such as #HitlerWasRight and #DeathToMuslim on social media platform X.
-[^2] https://www.nytimes.com/2023/11/15/technology/hate-speech-israel-gaza-internet.html Additionally, The Council on American-Islamic Relations reported receiving 774 help requests and bias reports from Muslims in the USA within a 16-day period.
-[^3] https://www.cair.com/press_releases/cair-reports-sharp-increase-in-complaints-reported-bias-incidents-since-107/ While digital interconnectivity facilitates swift information sharing, it simultaneously amplifies the spread and impact of HS, transcending geographical boundaries.
+The recent Israel–Hamas conflict has notably escalated both anti-Muslim and anti-Semitic sentiments worldwide, evidenced by the trending of hashtags such as #HitlerWasRight and #DeathToMuslim on the social media platform X.
+[^2] https://www.nytimes.com/2023/11/15/technology/hate-speech-israel-gaza-internet.html Moreover, the Council on American–Islamic Relations reported receiving 774 help requests and bias reports from Muslims in the USA within a 16-day period.
+[^3] https://www.cair.com/press_releases/cair-reports-sharp-increase-in-complaints-reported-bias-incidents-since-107/ While digital interconnectivity facilitates swift information sharing, it simultaneously amplifies the spread and the impact of HS, transcending geographical boundaries.
 Technological advancements have transformed the expression of HS, leading to its manifestation in various novel forms.
-Traditionally, HS was predominantly text-based, found in written materials Ana (1999), or verbalized in posts, broadcasts, and public speeches Nielsen (2002).
-However, the digital era has ushered in more complex and subtle variants of HS, engaging multiple sensory modalities.
-A notable instance is vision-language HS, which fuses visual elements with textual content, commonly disseminated through captioned images and memes Kiela et al.
-(2020); Fersini et al.
-(2022); Bhandari et al.
+Traditionally, HS was predominantly text-based, found in written materials Rini et al.
+(2020), or verbalized in posts, broadcasts, and public speeches Nielsen (2002).
+The digital era has ushered in more complex and subtle variants of HS, engaging multiple sensory modalities.
+A notable instance is vision-language HS, which fuses visual elements with text, commonly disseminated through captioned images and memes Uyheng et al.
+(2020); Kiela et al.
 (2020).
@@
 (2023).
-Figure 1 exemplifies various HS forms targeting immigrants, underscoring animosity towards individuals of diverse nationalities.
-The text-based approach overtly projects hostile attitudes towards immigrants in the host country.
+Figure 1 shows various HS forms targeting immigrants, underscoring animosity towards individuals of diverse nationalities.
+The text-based approach overtly projects hostile attitudes towards them in the host country.
 In vision-language HS, visual (e.g., a person preparing to shoot) and textual elements (e.g., sighting an illegal immigrant mowing the lawn) jointly convey antagonism.
 The figure also includes a music parody, integrating derogatory visuals with discriminatory audio lyrics, to showcase contempt for immigrants.
-While the digital landscape has seen increasing diversity in HS forms, existing research surveys Rini et al.
+While existing research surveys Rini et al.
 (2020); Chhabra and Vishwakarma (2023); Subramanian et al.
-(2023) predominantly concentrate on textual HS, often overlooking the complexity introduced by multimodal content.
-This survey aims to address this critical gap.
-Unlike prior surveys, we provide a comprehensive analysis of HS across various digital platforms, encompassing not only text but also visual, auditory, and combined multimodal expressions.
-We delve into the nuanced ways in which HS manifests in these different formats, offering insights into their unique characteristics and the challenges they pose for moderation.
-Furthermore, our survey underscores the pivotal role of large language models (LLMs) and large multimodal models (LMMs) in moderating HS.
-These advanced models, which can process and interpret multiple types of data simultaneously, are crucial in understanding and countering the multifaceted nature of modern HS.
-We critically assess the existing solutions, evaluating their effectiveness and highlighting areas for improvement.
-Our survey presents a novel perspective on HS moderation, shifting the focus from traditional text-based analyses to multimodal approaches.
-For this survey, we methodically examined research on moderating various types of hate speech, including text, images, videos, and audio.
-We searched with keywords like ‘hate speech’, ‘multimodal hate speech’, ‘harmful memes’, etc., on various scholarly platforms such as Google Scholar, DBLP, IEEE Xplore, and ACM Digital Library.
+(2023) have largely focused on text-based HS, they often overlook the complexity of multimodal content.
+Our survey addresses this gap by offering a comprehensive analysis of HS across various digital platforms, including text, visual, auditory, and combined multimodal expressions.
+We explore the distinct ways HS manifests in these formats, providing insights into their characteristics and moderation challenges.
+Additionally, we emphasize the critical role of large language models (LLMs) and large multimodal models (LMMs) in moderating HS, given their ability to process and interpret diverse data types.
+This survey critically evaluates existing solutions, identifies areas for improvement, and advocates for a shift towards multimodal approaches in HS moderation.
 In summary, our paper not only bridges the gap in the existing literature by providing a detailed exploration of multimodal HS but also paves the way for future research in this area.
 We aim to inspire advancements in HS moderation technology, particularly in the development and refinement of large models, which are imperative for tackling the complex and ever-changing nature of online HS.
+Paper Collection.
+We systematically examined research pertaining to the moderation of various types of hate speech, encompassing text, images, videos, and audio.
+Our search involved keywords such as ’hate speech’, ’multimodal hate speech’, ’hateful memes’, and similar terms, across scholarly platforms like Google Scholar, DBLP, IEEE Xplore, and ACM Digital Library.
+Among related research, we further selected state-of-the-art studies, with a particular interest in those using LLMs and LMMs.
+Due to the need for a manageable scope, we excluded works that did not leverage LLMs or LMMs, or focused narrowly on regional or multilingual aspects without broader relevance.
+This decision is not a reflection on the quality or importance of these works but rather a necessity to maintain a focused and coherent survey.
+Mod. Dataset Task Labels Source # Records
+Text WZ-LS Waseem and Hovy (2016) Det. [M.C.] Sexism, Racism, Neither Twitter 16,914
+GHC Kennedy et al. (2018) Det. [M.C.] VO, HD, CV [B] Implicitness [M.C.] Hate Targets Forums 27,665
+Stormfront de Gibert et al. (2018) Det. [B] Hateful StormFront 9,916
+DT Davidson et al. (2017) Det. [M.C.] Hateful, Offensive, Neither Twitter 24,802
+Founta Founta et al. (2018) Det. [M.C.] Offensive, Abusive, Hateful Speech, Aggressive, Cyberbullying, Spam, Normal Twitter 80,000
+DynaHate Vidgen et al. (2021) Det. [B] Hateful, [M.C.] Hate Targets, [M.C] Animosity, Derogation, Dehumanization, Threatening, Support H-M Adv 41,134
+SBIC Sap et al. (2020) Det. [B] Offensive [M.C] Hate Targets [B] Intent [B] Lewd [B] Group [B] In-Group Mixed 44,671
+Expl. [F.T.] Implied Statement
+IHC ElSherief et al. (2021) Det. [M.C.] Implicit, Explicit, Non-Hate [M.C.] Grievance, Incitement, Inferiority, Irony, Stereotypical, Threatening, Others Twitter 22,584
+Expl. [F.T.] Implied Statement
+HateXplain Mathew et al. (2021) Det. [M.C] Hate, Offensive, Normal [M.C.] Hate Targets Mixed 20,148
+Expl. [M.L] Text Rationales/Snippets
+NACL Masud et al. (2022) Det. [M.C.] Hate Intensity [M.L.] Hate Spans Mixed 4,423
+Ctr. [F.T.] Hate Speech Normalization
+CONAN Chung et al. (2019) Det. [M.C.] Hate Types [M.C.] Hate Sub-Topic Synthetic 14,988
+Ctr. [F.T.] CN Generation
+Multitarget CONAN Fanton et al. (2021) Det. [M.C.] Hate Targets GPT-2 5,003
+Ctr. [F.T.] CN Generation
+Counter Narratives Das et al. (2023) Ctr. [F.T.] CN Generation YouTube 9,119
+Img MMHS150K Gomez et al. (2020) Det. [B] Hateful Twitter 150,000
+FHM Kiela et al. (2020) Det. [B] Hateful Synthetic 10,000
+Finegrained FHM Mathias et al. (2021) Det. [B] Hateful [M.L.M.C] Protected Category [M.L.M.C] Protected Attacks Synthetic 10,000
+Misogynous Meme Det. [B] Misogynistic [B] Aggressive Mixed 800
+Gasparini et al. (2022) [B] Ironic
+MAMI Fersini et al. (2022) Det. [B] Misogyny [M.L.M.C.] Misogynous, Shaming, Stereotype, Objectification, Violence Mixed 10,000
+UA-RU Conflict Thapa et al. (2022) Det. [B] Hateful Twitter 5,680
+CrisisHateMM Bhandari et al. (2023) Det. [B] Hateful [B] Directed [M.C.] Hate Targets Mixed 4,723
+RUHate-MM Thapa et al. (2024) Det. [B] Hateful [M.C] Hate Targets Twitter 20,675
+HatReD Hee et al. (2023) Expl. [F.T] Explanations Synthetic 3,228
+Video Bangla Hate Videos Junaid et al. (2021) Det. [B] Hateful YouTube 300
+HateMM Das et al. (2023) Det. [B] Hateful [M.C.] Hate Targets Mixed 1,083
+MultiHateClip Wang et al. (2024) Det. [M.C] Hateful, Offensive, Normal YouTube & Bilibili 2,000
+Audio DeToxy Ghosh et al. (2021) Det. [B] Hateful Mixed 2M
+MuTox Costa-jussà et al. (2024) Det. [B] Hateful Mixed 116,000
+Table 1: Publicly available datasets for HS detection (Det.), HS explanation (Expl.) and counter HS (Ctr.).
+Abbreviation: M.L.: multi-label, M.C.: multi-class, M.L.M.C.: multi-label multi-class, B: binary, F.T: free-text, H-M Adv: Human-Machine Adversarial.
+Note that multilingual HS is out of the scope for the current review.
 2 Hate Speech
-Figure 2: Typology of HS based on modalities and tasks.
-The dark blue boxes are mature research areas with multiple studies; light grey boxes are ongoing research areas, while the hat can be found, and hatched boxes are unexplored topics.
+Figure 2: Typology of HS based on modalities and tasks. The dark blue boxes are mature areas with multiple studies; light grey boxes are ongoing research, and hatched boxes are unexplored topics.
 HS takes various forms — written text, images, spoken words, and multimedia content — each posing risks of violence, animosity, or prejudice against specific groups.
 This section reviews existing literature on HS, categorizing it into text-based, image-based, video-based, and audio-based types.
+For each HS form, we provide a detailed categorization across four tasks: detection, explanation, debiasing, and counter-speech.
+Detection identifies hateful content, forming the basis for further actions.
+Explanation promotes transparency by clarifying why content is flagged, building trust in automated systems.
+Debiasing is essential to refine detection systems, ensuring fairness and reducing bias.
+Counter-speech involves taking proactive steps to mitigate the impact of hate speech, fostering healthier online dialogue.
+Although these tasks address different aspects, they collectively form the foundation of an effective content moderation strategy, highlighting both the interconnectedness of HS forms and the research gaps in advancing multimodal HS moderation.
 Figure 2 illustrates the range of online HS forms.
-Additionally, Table 1 lists publicly accessible HS datasets for different modalities, providing researchers with essential resources for data collection and analysis, and supporting comprehensive and consistent research in HS detection and moderation.
+Additionally, Table 1 lists publicly accessible HS datasets in different modalities, providing researchers with essential resources for HS moderation.
 2.1 Text-based Hate Speech
-Text-based HS encompasses written or typed expressions manifesting across various digital formats, including social media posts Waseem and Hovy (2016); Founta et al.
+Text-based HS encompasses written or typed expressions manifested across online platforms, such as social media posts Waseem and Hovy (2016); Founta et al.
 (2018).
-Researchers addressing online HS have compiled datasets to detect and analyze its complexity and multifaceted nature, as detailed in existing literature.
-These studies explore diverse aspects of hateful and derogatory language, focusing on themes such as implicit HS ElSherief et al.
+Recent studies explored diverse aspects of hate and derogatory language, focusing on implicit HS Sap et al.
 (2020), targeted groups Kennedy et al.
+(2018); Yoder et al.
 (2022), and types of attacks ElSherief et al.
 (2021).
-As detection models evolve, it becomes imperative to understand and elucidate their decision-making processes, leading to the development of datasets for implicit statements and methodologies for bias analysis Sap et al.
-(2020); ElSherief et al.
+As HS detection models improve, it becomes imperative to understand and explain their decision-making processes, mitigating unintended bias Garg et al.
 (2023).
@@
 (2022).
-The detection of text-based HS is fraught with significant challenges.
-A primary obstacle is contextual understanding; HS often subtly or covertly conveys its message through sarcasm, irony, or cultural references.
-Linguistic variations, including slang, dialects, and unconventional language forms, exacerbate this intricacy, posing difficulties for automated recognition systems.
-Additionally, the multilingual nature of online content necessitates language-specific tools or models for effective detection across diverse linguistic contexts.
-A further complication is the rapid evolution of online language, which can quickly render existing detection models outdated.
+The detection of text-based HS poses numerous challenges.
+Detecting hate speech (HS) in a single statement often requires understanding dark humor and cultural nuances Hee et al.
+(2024).
+HS can express underlying intent through sarcasm, irony, or cultural references, which may not be immediately apparent.
+Linguistic variations, such as slang, dialects, and unconventional language use, further complicate the task.
+The challenge intensifies when considering the broader context of an utterance Nagar et al.
+(2023); Yu et al.
+(2022a), as statements that seem neutral in isolation may reveal hateful intent when viewed within a conversation.
+Conversely, what appears offensive might be harmless in context.
+Therefore, context-aware models are essential for accurately identifying HS by analyzing both individual statements and their surrounding situational context.
+Expanding the analysis to converssations, such as Reddit threads or WhatsApp chat, adds additional layers of complexity Naseem et al.
+(2019).
+The intent behind a single message can shift based on prior exchanges and the overall tone of the conversation.
+Furthermore, user-specific features may be important for HS detection Qian et al.
+(2018).
+Data such as a user’s posting history, profile, and behavior provide valuable context for identifying hate speech, though using such data raises ethical concerns, particularly regarding privacy.
 2.2 Image-based Hate Speech
 Image-based HS utilizes visual elements, such as photographs, cartoons, and illustrations, to propagate hate or discrimination against specific groups.
-A common manifestation of this type is memes, which typically consist of images combined with short overlaid text.
+A common manifestation of this HS form is memes, which typically consist of images combined with short overlaid text.
 Although memes often serve humorous or satirical purposes, they are increasingly used to spread hateful content online Kiela et al.
 (2020).
-The viral nature of memes on digital platforms further amplifies their impact, allowing them to reach vast audiences rapidly.
-To address the proliferation of hateful memes, recent studies have focused on developing datasets for their detection Kiela et al.
-(2020); Gomez et al.
-(2020) and for identifying specific targets and types of attacks within these memes Fersini et al.
-(2022).
-Beyond detection, new approaches are being pursued to analyze and mitigate bias in image-based HS detection models Hee et al.
+Recent studies have developed datasets for identifying HS Gomez et al.
+(2020), specific targets Mathias et al.
+(2021) and types of attacks Fersini et al.
+(2022) within these memes.
+Beyond detection, new approaches analyze and mitigate bias in image-based HS detection models Hee et al.
 (2023); Lin et al.
@@
 Additionally, new methodologies are emerging to counteract HS transmitted through memes Van and Wu (2023).
-Detecting image-based HS presents challenges due to the nuanced expression of offensive messages in visual content.
+Image-based HS presents new challenges due to the subtlety of offensive messages concealed within multiple modalities.
 Images, often embedding symbols, memes, or culturally specific visual cues, require deep cultural and contextual understanding for accurate interpretation.
-The combination of visual elements and text can subtly imply meanings not immediately evident Kiela et al.
+The visual elements and text can subtly imply meanings not immediately evident Kiela et al.
 (2020).
 For example, Figure 1 depicts a man with a gun and text suggesting hostility towards immigrants.
-Differentiating humor from hate in memes is particularly challenging, influenced by varying cultural, societal, and personal perspectives Schmid (2023).
+Differentiating humour from hate in memes is particularly challenging, influenced by varying cultural, societal, and personal perspectives Schmid (2023).
 2.3 Video-based Hate Speech
 Video-based HS presents a complex challenge, comprising a blend of visuals, audio tracks, and/or textual elements.
-This form of HS ranges from professionally produced propaganda to amateur videos on social media platforms like YouTube and TikTok.
+This form of HS ranges from professionally produced propaganda to amateur videos on social media platforms like YouTube and TikTok Das et al.
+(2023); Wang et al.
+(2024).
 The engaging nature of video content and its easy dissemination across digital networks significantly heighten its potential for harm.
 Echoing the concerns of image-based HS, video-based HS also contributes to the normalization of hateful ideologies and can profoundly influence public opinion.
-Contemporary research in this domain primarily focuses on identifying video-based HS and categorizing its various subtypes Wu and Bhandary (2020).
-Recent studies have explored detecting hateful segments within videos alongside multi-dimensional analysis.
-Nonetheless, the body of research on video-based HS is less developed than text-based and image-based HS, particularly in areas such as analyzing and mitigating model bias, elucidating decision-making processes, and devising counterstrategies.
+Contemporary research primarily focuses on identifying video-based HS and categorizing its various subtypes Wu and Bhandary (2020).
+Nonetheless, the amount of research on video-based HS is less developed than text-based and image-based HS, particularly in areas such as analyzing and mitigating model bias, elucidating decision-making processes, and devising counterstrategies.
 These gaps, likely stemming from the rapid pace of technological advancements and evolving digital trends, underscore the need for further research to promote a more harmonious online environment.
-Identifying HS within videos is a demanding and resource-intensive endeavor.
-This complexity stems from the amalgamation of various elements, such as text, images, and audio, within a single medium.
-Each component can independently harbor HS, thereby compounding the detection process.
+Detecting hate speech (HS) in videos is challenging and resource-intensive because it requires understanding various elements, including text, images, and audio, both independently and in combination.
+Each component can independently contain hateful content, further complicating the detection process.
 The duration of videos further exacerbates this challenge, as longer content necessitates more extensive review and analysis, with potential shifts in context over time.
 Moreover, subtle visual cues and sophisticated editing techniques can be employed to discreetly embed hate messages, making their detection by automated tools particularly challenging.
-Additionally, the analysis of video content requires considerable computational resources and time, posing a substantial challenge for organizations tasked with detecting and addressing HS in video formats.
+Additionally, video content analysis requires considerable computational resources and time, posing a substantial challenge for organizations to detect and address HS in video formats.
 2.4 Audio-based Hate Speech
 Audio-based HS entails the analysis of sound waves to discern elements such as pitch, intonation, and the contextual meaning of spoken words.
-This form of HS can originate from a variety of audio channels, including real-time conversations, podcasts, broadcasts, and other forms of audio media.
+This form of HS can originate from a variety of audio channels, including real-time conversations, podcasts, and other forms of audio media.
 The methodologies for addressing audio-based HS are diverse, targeting different facets of the issue.
 For instance, Barakat et al.
-Barakat et al.
 (2012) employed a straightforward keyword-based approach to identify segments of HS, while Wazir et al.
-Wazir et al.
 (2020) engaged in a detailed classification of offensive categories in audio-based HS, showcasing a nuanced method of understanding and categorizing this form of HS.
@@
 3 Methodology
-Mod. Dataset Task Labels Source # Records
-Text WZ-LS Waseem and Hovy (2016) Det.
-[M.C.] Sexism, Racism, Neither
-Twitter 16,914
-GHC Kennedy et al. (2018) Det.
-[M.C.] VO, HD, CV [B] Implicit, Explicit [M.C.] Hate Targets
-Forums 27,665
-Stormfront de Gibert et al. (2018) Det.
-[B] Hateful
-StormFront 9,916
-DT Davidson et al. (2017) Det.
-[M.C.] Hateful, Offensive, Neither
-Twitter 24,802
-HatEval Basile et al. (2019) Det.
-[B] Hateful [B] Target [B] Aggressive
-Twitter 19,600
-Founta Founta et al. (2018) Det.
-[M.C.] Offensive, Abusive, Hateful Speech, Aggressive, Cyberbullying, Spam, Normal
-Twitter 80,000
-SBIC Sap et al. (2020) Det.
-[B] Offensive [B] Intent [B] Lewd [B] Group [B] Hate Targets [B] In-Group
-Mixed 44,671
-Expl.
-[F.T.] Implied Statement
-IHC ElSherief et al. (2021) Det.
-[M.C.] Implicit, Explicit, Non-Hate [M.C.] Grievance, Incitement, Inferiority, Irony, Stereotypical, Threatening, Others
-Twitter 22,584
-Expl.
-[F.T.] Implied Statement
-HateXplain Mathew et al. (2021) Det.
-[M.C] Hate, Offensive, Normal [M.C.] Hate Targets
-Mixed 20,148
-Expl.
-[M.L] Text Rationales/Snippets
-DynaHate Vidgen et al. (2021) Det.
-[B] Hateful, [M.C] Animosity, Derogation, Dehumanization, Threatening, Support, [M.C.] Hate Targets
-H-M Adv 41,134
-ToxiGen Mathew et al. (2021) Det.
-[B] Toxic, Benign
-GPT-3 274,186
-NACL Masud et al. (2022) Det.
-[M.C.] Hate Intensity [M.L.] Hate Spans
-Mixed 4,423
-Ctr.
-[F.T.] Hate Speech Normalization
-CONAN Chung et al. (2019) Det.
-[M.C.] Hate Types [M.C.] Hate Sub-Topic
-Synthetic 14,988
-Ctr.
-[F.T.] CN Generation
-Multitarget CONAN Fanton et al. (2021) Det.
-[M.C.] Hate Targets
-GPT-2 5,003
-Ctr.
-[F.T.] CN Generation
-Counter Narratives Das et al. (2023) Ctr.
-[F.T.] CN Generation
-YouTube 9,119
-Img MMHS150K Gomez et al. (2020) Det.
-[B] Hateful
-Twitter 150,000
-FHM Kiela et al. (2020) Det.
-[B] Hateful
-Synthetic 10,000
-Finegrained FHM Mathias et al. (2021) Det.
-[B] Hateful [M.L.M.C] Protected Category [M.L.M.C] Protected Attacks
-Synthetic 10,000
-Misogynous Meme Det.
-[B] Misogynistic [B] Aggressive
-Mixed 800
-Gasparini et al. (2022)
-[B] Ironic
-MAMI Fersini et al. (2022) Det.
-[B] Misogyny [M.L.M.C.] Misogynous, Shaming, Stereotype, Objectification, Violence
-Mixed 10,000
-UA-RU Conflict Thapa et al. (2022) Det.
-[B] Hateful
-Twitter 5,680
-CrisisHateMM Bhandari et al. (2023) Det.
-[B] Hateful [B] Directed [M.C.] Hate Targets
-Mixed 4,723
-HatReD Hee et al. (2023) Expl.
-[F.T] Explanations
-Synthetic 3,228
-Video HateMM Das et al. (2023) Det.
-[B] Hateful [M.C.] Hate Targets
-Mixed 1,083
-Bangla Hate Videos Junaid et al. (2021) Det.
-[B] Hateful
-YouTube 300
-Audio DeToxy Ghosh et al. (2021) Det.
-[B] Hateful / Non-hateful
-Mixed 2M
-MuTox Costa-jussà et al. (2024) Det.
-[B] Hateful / Non-hateful
-Mixed 116,000
-Table 1: Publicly available datasets for HS detection (Det.), HS explanation (Expl.) and counter HS (Ctr.).
-Abbreviation: M.L.: multi-label, M.C.: multi-class, M.L.M.C.: multi-label multi-class, B: binary, F.T: free-text, H-M Adv: Human-Machine Adversarial.
-Note that multilingual HS is out of the scope for the current review.
-In this section, we review state-of-the-art methodologies that have made significant contributions to key areas of HS research: HS detection, HS explanation, HS debiasing, and counter speech.
-We emphasise research studies incorporating LLMs or LMMs, aiming to garner comprehensive insights into emerging trends within this domain.
-This allows us to understand how these advanced models enhance our understanding and management of HS in various forms.
+This section reviews the state-of-the-art methodologies that have significantly contributed to primary areas of HS research, particularly those involving large models.
+First, we discuss the recent capabilities of large models (Section 3.1).
+Subsequently, we explore studies in four important HS areas: detection (Section 3.2), explanation (Section 3.3), debiasing (Section 3.4), and counter-speech (Section 3.5), focusing on works using large models.
+This review highlights the emerging trends, providing insights into how large models can be used to understand and address HS in its various forms.
 3.1 Large Models
@@
 These foundation models have opened new avenues for identifying and mitigating hateful content, which requires nuanced understanding of language and context.
-In this survey, we regard both LLMs and LMMs as models with several billion parameters, aligning with the definition widely accepted and analyzed in numerous studies of large-scale models Luo et al.
-(2023).
+Here, we regard LLMs and LMMs as models with several billion parameters, aligning with the definition widely accepted and analyzed in numerous studies of large-scale models Luo et al. (2023).
 3.2 Hate Speech Detection
 The leading detection techniques for HS vary according to the modality of the content, encompassing approaches from transformer-based models to spectrogram-based classification models.
-For instance, AngryBERT Awal et al.
-(2021) employs a multi-task learning approach to fine-tune BERT specifically for binary text-based HS detection.
+For text-based HS detection, approaches range from embedding-based methods to advanced neural models Cao et al.
+(2020); Davidson et al.
+(2017); Badjatiya et al.
+(2017); Fortuna and Nunes (2018).
+AngryBERT Awal et al.
+(2021) fine-tunes BERT using a multi-task learning strategy for binary text HS detection.
 PromptHate Cao et al.
-(2022) integrates in-context learning with a demonstration sampling strategy for mask prompt fine-tuning on RoBERTa, targeting hateful memes.
-In audio-based HS classification, studies Boishakhi et al.
+(2022) combines demonstration sampling and in-context learning to fine-tune RoBERTa for hateful meme detection.
+In audio-based HS detection, ensemble techniques such as AdaBoost, Naive Bayes, and Random Forest have been employed.
+Boishakhi et al.
 (2021); Ibañez et al.
-(2021) have leveraged ensemble techniques involving AdaBoost, Naive Bayes, and random forest.
-Additionally, CNNs have been utilized to convert audio into spectrograms Medina et al.
-(2022), with self-attentive CNNs Yousefi and Emmanouilidou (2021) employed to extract audio features.
-For video-based HS classification, a combination of BERT, ViT, and MFCC has been used for text, image, and audio modality analysis, respectively Das et al.
 (2021).
-Note that video-based and audio-based HS detection represent emerging fields, offering substantial scope for further development and innovation.
-Transformer-based models have significantly advanced the detection of text-based and image-based HS; yet they encounter specific challenges.
+CNNs are also used to convert audio into spectrograms Medina et al.
+(2022), with self-attentive CNNs extracting audio features Yousefi and Emmanouilidou (2021).
+For video-based HS detection, a combination of BERT, ViT, and MFCC has been used for text, image, and audio modality analysis, respectively Das et al.
+(2023).
+Note that audio-based and video-based HS detection are emerging areas with significant potential for future advancements.
+Transformer-based models have significantly advanced the detection of text-based and image-based HS; yet they encounter challenges.
 For text-based models, a major hurdle is generalizing to out-of-distribution datasets, often hindered by limited vocabulary and the rarity of implicit HS in many datasets Ocampo et al.
@@
 Ocampo et al.
-Ocampo et al.
 (2023a) introduced a method using GPT-3 to generate implicit HS, aiming to both challenge and improve HS classifiers.
 Concurrently, Wang et al.
-Wang et al.
 (2023b) developed a technique for optimizing example selection for in-context learning in LLMs.
 In image-based HS, the primary challenge lies in deciphering implicit hate messages within memes.
-This difficulty often stems from the loss of information during the extraction of text-based features from images, a common step in many methodologies Lee et al.
+This often stems from the loss of information during the extraction of text-based features from images, a common step in many methodologies Lee et al.
 (2021); Pramanick et al.
@@
 To address these challenges, recent strategies include employing LMMs with prompting techniques and/or knowledge distillation.
-ProCap Cao et al.
+Pro-Cap Cao et al.
 (2023) addresses the issue of information loss in image-to-text conversion by prompting an LMM in a QA format, enhancing the generated caption’s quality and informativeness.
@@
 (2022).
-One proposed solution involves training supervised models that not only categorize HS but also provide rationales for these classifications Sap et al.
-(2020); ElSherief et al.
-(2021); Hee et al.
-(2023).
-Sap et al.
+One proposed solution involves training supervised models that not only categorize HS but also provide rationales for these classifications.
 Sap et al.
 (2020) and ElSherief et al.
-ElSherief et al.
 (2021) developed text-based HS datasets with human-annotated explanations, setting benchmarks for identifying underlying hate.
 Similarly, Hee et al.
-Hee et al.
 (2023) compiled a dataset for hateful memes, complete with human-annotated explanations and benchmarks.
-However, the process of gathering human-written explanations is not only time-consuming but also susceptible to individual biases.
+However, collecting human-written explanations is not only time-consuming but also susceptible to individual biases.
 Moreover, it involves the risk of subjecting human annotators to prolonged exposure to HS, which can have adverse psychological effects.
@@
 For instance, Wang et al.
-Wang et al.
 (2023a) demonstrated that GPT-3 can craft convincing and effective explanations for HS, a finding substantiated by extensive human evaluations.
 Additionally, HARE Yang et al.
-(2023a) introduces two innovative methods for generating rationales in HS detection, which have been shown to improve both the training process and the performance of detection models.
+(2023a) introduces two prompting methods that generate rationales for HS, enhancing the training of HS detection models and improving its performance.
 This approach presents an alternative means of developing insightful explanations, while simultaneously mitigating the risks associated with prolonged human exposure to HS.
@@
 Sap et al.
-Sap et al.
 (2019) found that two widely-used corpora exhibit bias against African American English, which increases the likelihood of classifying tweets in this dialect as hateful.
-Hee et al.
 Hee et al.
@@
 Beyond merely identifying biases, various studies have introduced innovative methods to reduce these biases within models.
-Kennedy Kennedy et al.
+Kennedy et al.
 (2020) developed a regularization technique utilizing SOC post-hoc explanations to address group identifier bias.
 Similarly, Rizzi et al.
-Rizzi et al.
 (2023) observed that models exhibit biases towards terms linked with stereotypical notions about women, such as dishwasher and broom.
 To counteract this, the authors proposed a bias mitigation strategy using Bayesian Optimization, which effectively lessened the bias while preserving overall model performance.
-These efforts underscore the critical importance of not only recognizing but actively mitigating bias in HS detection models.
+These efforts underscore the critical importance of not only recognizing, but also actively mitigating bias.
 This is especially vital as large models increasingly dominate the landscape for generating explanations and enabling transfer learning.
-Addressing biases within these large models represents a significant opportunity to advance new methodologies and strategies for both the identification and reduction of bias.
 3.5 Counter Speech
-The approach to counter HS emphasizes creating non-aggressive reactions, aimed at curbing the spread of HS and potentially converting it into respectful and inoffensive language.
-The primary approach for counter HS involves producing counter-narratives (CN), which rely on reliable evidence, logical arguments, and different perspectives to challenge the hate directed at marginalized groups Chung et al.
-(2021).
-Multitarget CONAN Fanton et al.
-(2021) uses GPT-2 to generate potential CN candidates and then engages human evaluators to identify effective CNs, leading to a dataset of 4,07840784,0784 , 078 CNs.
-An alternative approach involves either diminishing (i.e., normalization) or eliminating (i.e., correction) the level of hate in HS.
+The approach to countering HS focuses on generating non-aggressive responses that either reduce the spread of HS or transform it into respectful and inoffensive speech.
+Recent research categorizes counter-speech into various response types and emphasizes the importance of contextual understanding.
+Yu et al.
+(2023) developed a taxonomy of responses to HS, showcasing the diversity of counter-speech tactics.
+Mathew et al.
+(2019) proposed context-specific strategies such as narrative persuasion and active rebuttal.
+CONAN Chung et al.
+(2021) focused on generating counter-narratives that challenge hate directed at marginalized groups using reliable evidence, logical arguments, and diverse perspectives.
+These non-aggressive strategies reduce the spread of hate speech and foster positive discourse.
+Beyond generating non-aggressive responses, other approaches involve diminishing (i.e., normalization) or eliminating (i.e., correction) the level of hate in HS.
 NACL Masud et al.
-(2022) employs neural networks to paraphrase sections of text containing hate, effectively lessening the intensity of hate in HS.
-Van and Wu Van and Wu (2023) utilized LMM to identify and correct HS, substituting the text in memes with positive and respectful language.
-These studies underscore the critical role of generative models in annotating and developing counter speech strategies.
-This further signifies the future opportunities of LMMs and LMMs in enhancing approaches and techniques to combat hate speech effectively.
+(2022) used neural networks to paraphrase hate speech, effectively lowering the intensity of hate.
+Van and Wu (2023) prompted LMMs to correct HS in memes by replacing hateful text with positive and respectful language.
+These studies underscore the critical role of generative models in annotating and developing counter-speech strategies.
+This further signifies the future opportunities of LLMs and LMMs in enhancing approaches to combat hate speech.
 4 Challenges
@@
 Data Complexity, Quality, and Sourcing.
-Addressing multimodal HS brings to the fore complex challenges concerning data complexity, quality, and sourcing.
-In text and image-based contexts, issues range from reliance on pre-trained models to the nuanced interpretation of visual cues Cao et al.
-(2020); Awal et al.
-(2021); Sridhar and Yang (2022).
-For audio and video content, challenges are compounded by factors like varied accents, background noise, and the ambiguous nature of toxic content, coupled with inconsistent audio quality Yousefi and Emmanouilidou (2021).
+The subtlety of some hate speech, known as implicit hate speech, presents a considerable challenge in identifying and understanding the underlying intent, as these intents can hide within seemingly neutral language or actions Sap et al.
+(2020); Ocampo et al.
+(2023b); Kiela et al.
+(2020).
+This difficulty highlights the complexity of human communication and biases, where hateful messages can be conveyed indirectly or through coded language.
 Furthermore, sourcing data from diverse platforms such as Gab, YouTube, and 4chan introduces difficulties in standardization and interpretation Mariconti et al.
 (2019).
-Additionally, the uneven distribution of profanity across datasets poses significant obstacles for accurate model training.
-These challenges underscore the need for advanced methods capable of navigating the intricate and multifaceted nature of data across different modalities.
+Additionally, the uneven distribution of hate instances across datasets poses significant obstacles for accurate model training.
+Cao and Lee (2020).
+These challenges underscore the need for advanced methods capable of navigating the intricate and multifaceted nature of data.
 Model Performance and Generalizability.
 Recent research highlights the importance of enhancing HS detection models for adaptability in various scenarios and contexts.
-An exemplary example is making HS detection generalizable and effective across different domains Awal et al.
+An exemplary example is making HS detection generalizable and effective across domains Awal et al.
 (2021), underscoring the need for models to be versatile and not overly reliant on specific content cues such as domain, region, demography, and more.
 The development of systems like VulnerCheck Mariconti et al.
-(2019) exemplifies the demand for models that perform well regardless of the context, emphasizing the importance of creating models that can adapt to the ever-evolving nature of online material.
-Such adaptability is crucial for identifying and managing new hateful content, especially those designed to bypass advanced AI technologies.
-The adoption of technologies, like Few-Shot Learner (FSL), for quick adaptation to this evolving landscape shows a promising direction [^4] https://ai.meta.com/blog/harmful-content-can-evolve-quickly-our-new-ai-system-adapts-to-tackle-it/.
-However, it is imperative that these technologies not only understand the content but also integrate critical aspects of cultural, behavioral, and conversational contexts into their analysis.
-Modality and Expression Variabilities.
-Research has highlighted the complexities involved in interpreting expressions of HS across individual and combination of various modalities, particularly in cyberbullying scenarios Boishakhi et al.
+(2019) exemplifies the demand for models that perform well regardless of the context, and that can adapt to the ever-evolving nature of online material.
+Such adaptability is crucial for identifying and managing new hateful content, especially such designed to bypass advanced AI technologies.
+The adoption of technologies, like Few-Shot Learner (FSL), for quick adaptation to this evolving landscape is a promising direction.
+[^4] https://ai.meta.com/blog/harmful-content-can-evolve-quickly-our-new-ai-system-adapts-to-tackle-it/ However, it is imperative that these technologies not only understand the content, but also integrate critical aspects of cultural, behavioral, and conversational contexts.
+Expression and Modality Variabilities.
+Research has highlighted the complexities of interpreting hate speech (HS) across various modalities Boishakhi et al.
+(2021); Kiela et al.
 (2020).
-The main challenges in this area include the limitations of pre-trained models in accurately capturing the subtleties of HS expressed via multiple modalities Cao and Lee (2020), the difficulty in balancing sensitivity and specificity in detection algorithms Cao et al.
-(2020), and the imperative for models to distinguish between explicit and implicit forms of HS Kim et al.
+In text-based HS, implicit hate messages often use dark humor or sarcasm to obscure their true intent, making detection particularly challenging Sap et al.
+(2020); ElSherief et al.
 (2021).
-Additionally, the potential for misinterpretation and the sensitivity of models to specific trigger language Sridhar and Yang (2022) pose significant challenges.
-Addressing these issues is crucial and requires concerted efforts in future research endeavors.
+For image-based HS, models face difficulty identifying subtle cross-modality nuances that convey the underlying message Hee et al.
+(2022); Rizzi et al.
+(2023).
+These challenges become even more pronounced with audio and video content due to factors such as accents, background noise, inconsistent audio quality, and the inherently ambiguous nature of toxic content Yousefi and Emmanouilidou (2021).
+Additionally, the potential for misinterpretation and the models’ sensitivity to specific trigger words exacerbate these issues Sridhar and Yang (2022).
+Addressing these variabilities is essential and calls for dedicated efforts in future research.
 Contextualization.
-Recent research underscores the complexities of identifying and countering online hate speech (HS), requiring an accurate and comprehensive understanding of HS, ranging from the targeted victims to its implicit meanings Sap et al.
-(2020); Masud et al.
+Effective hate speech moderation requires a nuanced understanding of the HS context.
+Multi-turn interactions in social media conversations, like Reddit threads or Twitter discussions, play a key role in detecting implicit hate speech Ghosh et al.
+(2023) and generating counter speech Yu et al.
+(2022b).
+Additionally, Meng et al.
+proposed DRAG++, a model that predicts hate intensity by analyzing both the content and the full context of conversation threads Meng et al.
 (2023).
-An effective strategy to counter hate speech depends on this precise interpretation, enabling the creation of context-specific responses.
-The main obstacles in accurately decoding hate speech involve recognizing the subtle nuances in language use, such as wordplays, dual meanings, or ambiguous expressions Kennedy et al.
+Furthermore, geographic-specific factors, such as local slang and cultural differences, influence a model’s ability to generalize across different regions Lee et al.
 (2023).
-Additionally, the interpretation of hate speech often depends on its context, including cultural, societal, and situational elements Rizzi et al.
-(2023).
-These challenges highlight the need for sophisticated algorithms capable of interpreting language within its contextual usage, thereby enhancing the accuracy and effectiveness of hate speech detection and mitigation strategies.
+These challenges highlight the need for sophisticated algorithms capable of interpreting language within its contextual usage, thereby enhancing the accuracy and effectiveness of hate speech moderation strategies.
 Emerging Domains.
@@
 (2022).
-The core of these challenges lies the need to adapt current hate speech (HS) detection methods to new contexts and to develop new strategies specifically designed for the unique characteristics of these platforms.
+The core of these challenges lies the need to adapt current HS detection methods to new contexts and to develop new strategies specifically designed for the unique characteristics of these platforms.
 The dynamic and immersive nature of these emerging environments necessitates a re-evaluation and potential re-engineering of current HS detection and mitigation strategies.
 Future research requires a deep understanding of both technological advancements and the social dynamics within these virtual spaces to ensure effectiveness in detecting and mitigating HS within these evolving digital landscapes.
-Bias and Ethical Concerns:
-Confronting bias and upholding ethical considerations in AI systems, particularly in the context of HS detection, represents a challenge, as detailed in works such as Sandulescu (2020).
-These issues are not just technical but also moral, encompassing the need to ensure that AI systems operate equitably and do not inadvertently perpetuate or amplify existing societal biases.
-The development and implementation of responsible AI systems, therefore, require a multidisciplinary approach that integrates technical proficiency with ethical and societal awareness.
-This necessity underlines the importance of continuous efforts to refine AI algorithms, ensuring they align with ethical standards and societal values.
+Bias and Ethical Concerns.
+Addressing bias and upholding ethical considerations in HS detection systems poses a significant challenge.
+Several studies have highlighted these concerns, introducing functional tests for evaluating HS detection models Röttger et al.
+(2022); Ng et al.
+(2024); Röttger et al.
+(2020).
+These challenges are not purely technical but also moral, underscoring the importance of ensuring HS systems operate equitably and do not perpetuate societal biases.
+Developing responsible HS systems, therefore, require a multidisciplinary approach that combines technical expertise with ethical and societal awareness, ensuring alignment with ethical standards and societal values.
 In summary, the areas of HS detection and moderation are confronted with multifaceted challenges.
@@
 The future of this field hinges on our ability to continuously evolve and innovate, ensuring that our approaches remain relevant, effective, and ethically sound in an ever-changing digital landscape.
-5 Future Directions: Role of Large Models
-To conclude this survey, we explore prospective research opportunities in utilizing large-scale models for moderating hate content.
+5 Future Directions
 Cross-Modality Context Understanding.
-As hate speech extends beyond mere text to encompass multiple forms of media (multimodality), it becomes crucial for models to possess a proficient understanding of context across these different modalities.
-Hence, it is imperative that models not only identify hateful content within text or images separately but also grasp how the combination of text and images can alter the overall message Kiela et al.
+As hate speech extends beyond mere text to encompass multiple forms of media (multimodality), it becomes crucial for models to have a proficient understanding of context across modalities.
+Hence, it is imperative that models not only identify hateful content within text or images separately, but also grasp how the combination of text and images can alter the message Kiela et al.
 (2020).
 For instance, an image that is benign on its own might become hateful when paired with specific text.
-Research could focus on developing models that more effectively understand context across different modalities.
+Research could focus on developing models that more effectively understand context across modalities.
 Low-Resource Hate Speech Adaptation.
-Domain adaptation between related tasks has gained significant attention, with transfer learning emerging as a prominent strategy for addressing this challenge.
+Domain adaptation between related tasks has gained significant attention.
 In the domain of hate speech, an exemplary application is the cross-lingual transfer learning for detecting hate speech across different languages.
@@
 Nguyen et al.
-Nguyen et al.
 (2023) demonstrated how providing cultural common-sense knowledge can alter GPT-3’s behaviour, leading it to produce more accurate and culturally sensitive questions.
@@
 6 Conclusion
-This survey has highlighted significant advancements in HS moderation, underscoring the pivotal role of large language and multimodal models.
+We highlighted the advancements in HS moderation, underscoring the pivotal role of LLMs and LMMs.
 Despite these strides, challenges remain, particularly in inclusivity and nuanced detection.
 Future research should focus on developing AI methodologies that are more context-aware and ethically governed.
-This endeavor is not only a technological challenge but also a moral imperative, necessitating interdisciplinary collaboration.
-As we advance, it is crucial to balance innovation with responsibility, striving for a digital landscape that is safer and more inclusive for all.
+This endeavor is not only a technological challenge, but also a moral imperative, necessitating interdisciplinary collaboration.
+As we advance, it is crucial to ensure that technological advancements are matched with a commitment to responsibility, striving for a digital environment that is secure and welcoming for everyone.
+Acknowledgement
+This research is supported by the Ministry of Education, Singapore, under its Academic Research Fund Tier 2 (Award ID: MOE-T2EP20222-0010).
+Any opinions, findings and conclusions or recommendations expressed in this material are those of the authors and do not reflect the views of the Ministry of Education, Singapore.
+Tanmoy Chakraborty acknowledges Anusandhan National Research Foundation (CRG/2023/001351) for the financial support.
+Limitations
+There are two limitations to the scope and coverage of our survey paper.
+Scope.
+In this survey paper, we specifically focus on the role of large models in multimodal “hateful speech moderation”.
+We recognize that there is extensive research on toxic and harmful content that is closely related to hatefulness.
+However, hate speech has a distinct definition that attacks need to discriminate against a group of people based on specific traits such as race, gender and sexual orientation.
+Hence, while these closely related areas are significant, they fall outside the scope of this survey paper.
+We also recognize different forms of hate speech can exist in multiple languages, resulting in exciting research on multilingual hate speech.
+However, the primary goal of this paper is to highlight the evolution and adaptation of hate speech in various forms of digital content.
+Therefore, the topic of multilingualism is beyond the scope of this paper.
+Research Paper Coverage.
+Although numerous research studies on hate speech have been conducted over the past decades, we have focused on state-of-the-art works that either employed large models in their studies or pioneered specific hate speech tasks.
+This selection enables us to maintain the brevity of this survey paper while focusing our discussion on the promising areas of using large models for hate speech moderation.